OpenStudy (anonymous):

why is 18^1/5 called the 5th root of 18?

OpenStudy (anonymous):

by definition

OpenStudy (jdoe0001):

\(\Large \bf a^{\frac{{\color{blue} n}}{{\color{red} m}}} = \sqrt[{\color{red} m}]{a^{\color{blue} n}}\qquad thus\quad 18^{\frac{{\color{blue}{ 1}}}{{\color{red}{ 5}}}}\implies ?\)

OpenStudy (anonymous):

@satellite73 i dont understand what you mean by that..

OpenStudy (anonymous):

there is no such thing as multiplying a number by itself one fifth time we define \[\huge b^{\frac{1}{n}}\] to be \[\huge \sqrt[n]{b}\] in order to make the laws of exponents hold for fractional exponents as well as whole number exponents

OpenStudy (anonymous):

one law of exponents is \[\large (b^n)^m=b^{m\times n}\] that means if \(b^{\frac{1}{n}}\) is to make sense then since \[\large (b^{\frac{1}{n}})^n=b^1=b\] making \[\large b^{\frac{1}{n}}=\sqrt[n]{b}\]

OpenStudy (anonymous):

oh got it thanks for the help @satellite73

OpenStudy (anonymous):

yw